calculate the total no. of electrons in 18ml of water

 calculate the  total no. of electrons in 18ml of water 
 

Grade:12th pass

3 Answers

Bhavya
7 years ago
Dear student
Here we will use the the concept of density of water = 1g/cc
D=M/V
1 g/ml = M/18 ml
M= 18 g
moles of water = mass/molar mass = 18/18 = 1
1 mole of water contains = NA molecules
total electrons in water = 2(H) + 8(O) = 10
1 molecule of water contains = 10 electrons
1 mole water contains = 10 * NA electrons
NA = Avogadro’s number
Laxit Saini
6 years ago
Density of water= 1gm/cm³ 1g/ml=18gm/18mlSo, 1mole/18gmNow 1molecule of H2O will have 10 electrons.....1+1+8Therefore 1 mole of H2O will have 10xNA electrons....so ans is 10NA
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your answer below.
 
Here we will use the the concept of density of water = 1g/cc
D=M/V
1 g/ml = M/18 ml
M= 18 g
moles of water = mass/molar mass = 18/18 = 1
1 mole of water contains = NA molecules
total electrons in water = 2(H) + 8(O) = 10
1 molecule of water contains = 10 electrons
1 mole water contains = 10 * NA electrons {NA = Avogadro’s number}
 
Hope it helps.
Thanks and regards,
Kushagra